Please help! Given the Vertex, Axis of Sym., and y-intercept what is the equation? Thanks!
vertex (4,11)
A of S = x=4
y-intercept (0,3)

Answer by scott8148(6628)

the vertex form of a parabolic (quadratic) equation is y-k=a(x-h)^2
__ vertex is (h,k) and a is a measure of the width of the opening of the parabola

using the vertex __ y-11=a(x-4)^2

substituting the y-intercept __ 3-11=a(0-4)^2 __ -8=a(16) __ -1/2=a

y-11=(-1/2)(x-4)^2